# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

### v1

@@ 6.2 Classification Task: Non-Order-Preserving Sequences of Partitions
Task
In our second experiment, we classify whether a sequence of partitions is order-preserving or not, i.e., whether a sequence $\theta$ is compatible with a total ordering on the set $X$. This task is of relevance in several areas, from the study of preference relations in utility theory in social sciences [Roberts, 2009] to the analysis of weak orderings and partition refinement algorithms in computer science [Habib et al., 1999].
Definition 36 (Order-preserving sequence of partitions).
When a partition $\theta(t_{m})$ is equipped with a total order $_{m}$ on the clusters it is called an ordered partition. [^2] The ranking $\tau_{m}:V_{m}\rightarrow\{1,\dots,|V_{m}|\}$ of the vertices $V_{m}$ in the Sankey diagram $S(\theta)$ is one example of a total order $_{m}$ on the clusters, see Section 3.1. Such a partition induces a total preorder $\lesssim_{m}$ on $X$ [Stanley, 2011], i.e., if $x]_{t}<_{m}[y]_{t}$ then $x\lesssim_{m}y$. We call the sequence of partitions $\theta$ order-preserving if there exist total orders $(<_{1},\dots,<_{M})$ such that the total preorders $(\lesssim_{1},\dots,\lesssim_{M})$ are compatible across the sequence, i.e., $\forall\ell,m$ we have $x\lesssim_{\ell}y$ iff $x\lesssim_{m}y$, $\forall x,y\in X$.
Figure 4: Difference between order-preserving ($y=0$) and non-order-preserving ($y=1$) sequences (**** indicates $p<0.0001$, Mann-Whitney U test).
Data
We carry out this classification task on synthetic data for which we have a ground truth. From the space of coarse-graining sequences of partitions $\Pi_{N}^{M}$, introduced in Definition 35, we generate a balanced dataset of 3,700 partitions $\theta\in\Pi_{500}^{30}$, half of which are order-preserving ($y=0$) and the other half are non-order-preserving ($y=1$). The loss of order-preservation is induced by introducing random swaps in the node labels across layers. See Appendix F.2 for details. For each of the generated $\theta$ we compute $\mathrm{CE}$, $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the computationally advantageous nerve-based MCbiF. We choose $N=500$ and $M=30$ to demonstrate the scalability of our method.
Table 2: Classification task. Test accuracy of logistic regression trained on different features.
@@
encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{CE}$
0.53 0.56 0.97 0.50
Results
Firstly, whereas we find no significant difference between the information-theoretical $\mathrm{CI}$ of order-preserving ($y=0$) and non-order-preserving ($y=1$) sequences, we observe a statistically significant increase of $\bar{c}_{0}$ and $\bar{c}_{1}$ for order-preserving sequences (Fig. 4).
For the classification task, we split our data into training (80%) and test (20%). For each feature map, we then train a logistic regression on the training split, and evaluate the accuracy on the test split, see Appendix F.2. We find that $\mathrm{HF}_{1}$ predicts the label $y=\{0,1\}$ encoding the (lack of) order preservation with high accuracy (0.95). In contrast, $\mathrm{CE}$ and the raw label encoding of $\theta$ cannot improve on a random classifier (Table 2). Our results thus demonstrate the high sensitivity of MCbiF to order-preservation in $\theta$ because non-order-preserving sequences induce 1-conflicts that we capture with $\mathrm{HF}_{1}$.
@@ 6.3 Application to Real-World Temporal Data
In our final experiment, we apply MCbiF to temporal sequences of partitions computed from real-world contact data of free-ranging house mice that capture the changes in the social network structure of the rodents over time [Bovet et al., 2022].
Data
@@
Each partition $\theta_{\tau}(t)$ describes mice social groupings for $N=281$ individual mice at weeks $t\in[1,\dots,9]$, i.e., the nine weeks in the study period (28 February-1 May 2017). Hence, each sequence captures the fine-graining of social groups over the transition from winter to spring. Each partition sequence is computed at temporal resolution $\tau>0$, where the parameter $\tau$ modulates how fine the temporal community structure is (Fig. 13). See Bovet et al. [2022] for details.
We use MCbiF to compare the temporal sequences $\theta_{\tau_{i}}$ for nine parameters $\tau_{i}$, $i=1,\dots,9$, as provided in Bovet et al. [2022]. See Appendix F.3 for details. For each of the nine partition sequences $\theta_{\tau_{i}}$, we compute $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the computationally advantageous nerve-based MCbiF, which induces a 50-fold reduction in computation time due to a much lower number of simplices (260 simplices for the nerve-based MCbiF instead of 116,700 for the original MCbiF).
Figure 5: (a) Analysis of non-hierarchical sequences of partitions $\theta_{\tau_{i}}$ compiled from the temporal social interactions of a mice population over a period of 9 weeks. Each $\theta_{\tau_{i}}$ is formed by a sequence of social groupings $\theta_{\tau_{i}}(t)$ over week $t$. Different sequences of partitions were computed as a function of the parameter $\tau_{i}$. We display Sankey diagrams and MCbiF feature maps for $\theta_{\tau_{i}}$ at three parameters $\tau_{i}$ ($i=2,4,8$) identified as robust in the original work by Bovet et al. [2022]. These three sequences $\theta_{\tau_{i}}$ exhibit different types of non-hierarchy, as shown by our topological feature maps and our measures of average 0-conflict ($\bar{c}_{0}$) and average 1-conflict ($\bar{c}_{1}$). (b) The $\theta_{\tau_{i}}$ ($i=2,4,8$) found in Bovet et al. [2022] as robust behaviours correspond to distinct topological characteristics of the sequences of partitions, as captured by the block structure in the distance between MCbiF Hilbert functions.
Results
@@
Bovet et al. [2022] identified that the temporal resolutions $\tau_{2}=1\,\mathrm{s}$, $\tau_{4}=60\,\mathrm{s}$ and $\tau_{8}=24\,\mathrm{h}$ lead to robust sequences of partitions. Using the Hilbert distance, i.e., the $L_{2}$-norm on the 0- and 1-dimensional MCbiF Hilbert functions, we find these temporal resolutions to be representative for three distinct temporal regimes characterised by different degrees of non-hierarchy, as measured by $\bar{c}_{0}$ and $\bar{c}_{1}$ (Fig. 5). In particular, high $\bar{c}_{0}$ indicates that mice tend to split off groups over time, and high $\bar{c}_{1}$ indicates that mice meet in overlapping subgroups but never jointly in one nest box. Note that $\theta_{\tau_{2}}$ has a strong non-hierarchical structure because the large-scale mice social clusters get disrupted in the transition to spring. In contrast, $\theta_{\tau_{8}}$ is more hierarchical as it captures the underlying stable social groups revealed by the higher temporal resolution. However, $\theta_{\tau_{4}}$ has the strongest hierarchy as indicated by a lower $\bar{c}_{0}$ and an absence of 1-conflicts ($\bar{c}_{1}=0$) and thus corresponds to a sweet spot in hierarchical organisation between the low and high temporal resolutions. Finally, the Hilbert distances also capture an increased time reversibility in the sequence $\theta_{\tau_{8}}$ due to the increased stability of social groupings over time, see Fig. 14 in Appendix F.3.
@@ 7 Conclusion
We have introduced the MCbiF, a novel bifiltration that encodes the cluster intersection patterns in a multiscale, non-hierarchical sequence of partitions, $\theta$. Its stable Hilbert functions $\mathrm{HF}_{k}$ quantify the topological autocorrelation of $\theta$ and measure non-hierarchy in two complementary ways: the Hilbert function at dimension $k=0$ captures the absence of a maximum with respect to the refinement order (0-conflicts), whereas the Hilbert function at dimension $k=1$ captures the emergence of higher-order cluster inconsistencies (1-conflicts). This is summarised by the measures of average 0-conflict $\bar{c}_{0}(\theta)$ and average 1-conflict $\bar{c}_{1}(\theta)$, which are global, history-dependent and sensitive to the ordering of the partitions in $\theta$. The MCbiF extends the 1-parameter MCF defined by Schindler and Barahona [2025] to a 2-parameter filtration, leading to richer algebraic invariants that describe the full topological information in $\theta$. We remark that the MCbiF is independent of the chosen clustering algorithm and can be applied to any (non-hierarchical) sequence of partitions $\theta$.
We demonstrate with numerical experiments that the MCbiF Hilbert functions provide topological feature maps that can be used for downstream machine learning tasks, and are shown to outperform information-based features on regression and classification tasks on non-hierarchical sequences of partitions. Moreover, the grounding of MCbiF features in algebraic topology enhances interpretability, a crucial attribute for XAI and applications to real-world data.
@@
Our analysis of the MCbiF MPH is restricted to dimensions 0 and 1 due to current limitations of the RIVET software [Lesnick and Wright, 2015] used in our numerical experiments. Analysing topological autocorrelation for higher dimensions would allow us to capture more complex higher-order cluster inconsistencies and could be the object of future research.
Furthermore, we focused here on Hilbert functions as our topological invariants because of their computational efficiency and analytical simplicity, which facilitates our theoretical analysis. In future work, we plan to use richer feature maps by exploiting the block decomposition of the MCbiF persistence module, which leads to barcodes [Bjerkevik, 2021], or by using multiparameter persistence landscapes [Vipond, 2020].
Another future direction is to use MCbiF to evaluate the consistency of assignments in consensus clustering [Strehl and Ghosh, 2002, Vega-Pons and Ruiz-Shulcloper, 2011]. Indeed, it can be shown that the values of the Hilbert function $\mathrm{HF}_{k}(s,t)$ that are further away from the diagonal ($s=t$) are more robust to permutations of the ordering of partitions in $\theta$ (see Proposition 41 in Appendix A.4), and, in particular, $\mathrm{HF}_{k}(t_{1},t_{M})$ only depends on the set of distinct partitions in the sequence $\theta([t_{1},\infty))$ and is independent to any permutation in their order. Hence, in future work, $\mathrm{HF}_{k}(t_{1},t_{M})$ could be used as an overall measure of consistency in $\theta$ in the context of consensus clustering.
Finally, we plan to analyse minimal cycle representatives of the MPH [Li et al., 2021] to localise 1-conflicts in the sequence of partitions, which is of interest to compute conflict-resolving partitions in consensus clustering, or to identify inconsistent assignments in temporal clustering [Liechti and Bonhoeffer, 2020].
@@
Detailed proofs of all theoretical results can be found in Appendix A. Extensive documentation of our experiments is presented in Appendix F. The dataset studied in Section 6.3 is publicly available at: https://dataverse.harvard.edu/file.xhtml?fileId=5657692. Python code for the MCbiF method is publicly available at: https://github.com/barahona-research-group/MCF
@@ Acknowledgments
JS acknowledges support from the EPSRC (PhD studentship through the Department of Mathematics at Imperial College London). MB acknowledges support from EPSRC grant EP/N014529/1 supporting the EPSRC Centre for Mathematics of Precision Healthcare. We thank Kevin Michalewicz, Asem Alaa and Christian Schindler for valuable discussions on the computational aspects of this project. We thank Alex Bovet for discussions about the temporal dataset studied in this paper. We also thank Arne Wolf for discussions about the stability of multiparameter persistence modules. This work benefited from discussions at the London-Oxford TDA seminar in November 2024.
@@
Appendix A Proofs of Theoretical Results
A.1 Proofs for Section 1
@@ A.2 Proofs for Section 4
Next, we provide formal definitions for algebraic properties of persistence modules, see Botnan and Lesnick [2023] for details.
@@
Let $\mathbf{Vect}$ denote the category of $\boldsymbol{k}$-vector spaces for a fixed field $\boldsymbol{k}$. For a partially ordered set $P$, a $P$-indexed persistence module is a functor $F:P\mapsto\mathbf{Vect}$. We say that:
@@
    $F$ is called pointwise finite-dimensional if $\dim(F_{\boldsymbol{a}})<\infty$ for all $\boldsymbol{a}\in P$.
@@
    $F$ is called block-decomposable if it decomposes into blocks $F\bigoplus_{B\in\mathcal{B}(F)}\boldsymbol{k}_{B}$ where $\mathcal{B}(F)$ is a multiset of blocks that depends on $F$.
We can now provide the proof for Proposition 6, which shows that the MCbiF persistence module (see Fig. 7) is pointwise finite-dimensional, finitely presented and block-decomposable. The proof relies on the equivalent nerve-based construction of the MCbiF (see Proposition 40), and the exactness of the persistence module from which block-decomposability follows [Cochoy and Oudot, 2020].
@@
To prove block-decomposability, we use the nerve-based MCbiF $(\tilde{K}^{s,t})_{t_{1}\leq s\leq t}$, which leads to the same persistence module, see Proposition 40. As the module is uniquely defined by its values on a finite grid, we can use Theorem 9.6 by Cochoy and Oudot [2020] that implies block-decomposability if the persistence module is exact. Hence, it suffices to show that for all $t_{1}\leq t\leq t^{\prime}\leq t^{\prime\prime}\leq t^{\prime\prime\prime}$ the diagram
@@
By construction of the MCbiF, $\tilde{K}^{t,t^{\prime\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. Furthermore, $\tilde{K}^{t,t^{\prime\prime}}=\tilde{K}^{t,t^{\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime}}$ and $\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}=\tilde{K}^{t^{\prime},t^{\prime\prime}}\cup\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}$. Without loss of generality, $_{k}, t=tt^{\prime}=t_{\ell}$, $^{\prime\prime}=t_{m}, t′′′=t$$^{\prime\prime\prime}=t_{n}$ for change points $t_{k}<t_{\ell}<t_{m}<t_{n}$ of $\theta$ such that $A(k,\ell)\cap A(m,n)=\emptyset$. Hence, $\tilde{K}^{t,t^{\prime}}\cap\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}=\emptyset$ and $\tilde{K}^{t^{\prime},t^{\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cap\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. This means that Eq. (12) is a Mayer-Vietoris sequence for all $k\geq 0$, implying exactness [Hatcher, 2002, p. 149] and proving the block decomposability [Cochoy and Oudot, 2020, Theorem 9.6]. ∎
@@
A.2.1 Proofs for Section 4.1
@@
(i) Moreover, $\nexists r\in[s,t]$ such that $x\sim_{r}\sim_{r}y$. In particular, $\nexists r\in[s,t]$ such that $\nexists C^{\prime\prime}\in\theta(r)$ with $C\subseteq C^{\prime\prime}$ and $C\subseteq C^{\prime\prime}$. Hence, $\nexists r\in[s,t]$ such that $\theta(r_{1})\leq\theta(r)$ and $\theta(r_{2})\leq\theta(r)$, implying that the subposet $\theta([s,t])$ has no maximum. This shows that every triangle 0-conflict is also a 0-conflict, proving statement (i). Note that the opposite is not true as illustrated by the example in Fig. 3b. ∎

### v2

@@ 5.2 Classification Task: Non-Order-Preserving Sequences of Partitions
In our second task, we classify whether a sequence of partitions is order-preserving, i.e., whether $\theta$ is compatible with a total ordering on the set $X$, see Definition 41 in Appendix D.2. This task is relevant in socio-economics (preference relations in utility theory in social sciences (Roberts, 2009)) and computer science (weak ordering and partition refinement algorithms (Habib et al., 1999)).
We carry out this classification task on synthetic data for which we have a ground truth. From the space of coarse-graining sequences of partitions $\Pi_{N}^{M}$ introduced in Definition 40, we generate a balanced dataset of 3,700 partitions $\theta\in\Pi_{500}^{30}$, half of which are order-preserving ($y=0$) and the other half are non-order-preserving ($y=1$). The loss of order-preservation is induced by introducing random swaps in the cluster assignments within layers. See Appendix D.2 for details. For each of the generated $\theta$ we compute $\mathrm{CE}$, $\mathrm{ARI}$, $\mathrm{MOD}$, $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the computationally advantageous nerve-based MCbiF. We choose $N=500$ and $M=30$ to demonstrate the scalability of our method.
Table 2: Classification task. Test accuracy of logistic regression trained on different features. See Appendix D.2 for confidence intervals.
@@
encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{CE}$ $\mathrm{ARI}$ $\mathrm{MOD}$
0.53 0.56 0.97 0.50 0.49 0.46
Whereas we find no significant difference between the baseline consensus indices of order-preserving ($y=0$) and non-order-preserving ($y=1$) sequences, we observe a significant increase of $\bar{c}_{0}$ and $\bar{c}_{1}$ for order-preserving sequences (Fig. 12 in Appendix D.2). For the classification task, we split our data into training (80%) and test (20%). For each feature map, we train a logistic regression on the training split and evaluate the accuracy on the test split, see Appendix D.2. We find that $\mathrm{HF}_{1}$ predicts the label $y=\{0,1\}$ encoding the (lack of) order-preservation with high accuracy (0.97). In contrast, $\mathrm{CE}$, $\mathrm{ARI}$, $\mathrm{MOD}$ and the raw label encoding of $\theta$ do not improve on a random classifier (Table 2). Our performance variability analysis (bootstrapped 95% confidence intervals) confirms that the model trained on $\mathrm{HF}_{1}$ consistently outperforms all models (Table 5 in Appendix D.2). This demonstrates the high sensitivity of MCbiF to order-preservation in $\theta$, due to non-order-preserving sequences inducing 1-conflicts that are captured by $\mathrm{HF}_{1}$.
Figure 3: (a) Analysing non-hierarchical sequences of temporal partitions $\theta_{\tau_{i}}$ compiled from social interactions of a mice population over 9 weeks. Each $\theta_{\tau_{i}}$ is a sequence of social groupings $\theta_{\tau_{i}}(t)$ observed over week $t$. Different sequences of partitions are computed as a function of the parameter $\tau_{i}$. Sankey diagrams and MCbiF feature maps are shown for $\theta_{\tau_{i}}$ at $\tau_{i}$ ($i=2,4,8$), identified as robust in Bovet et al. (2022). These three sequences exhibit different types of non-hierarchy, as shown by our topological feature maps and measures of average 0-conflict ($\bar{c}_{0}$) and average 1-conflict ($\bar{c}_{1}$). (b) The robust $\theta_{\tau_{i}}$ ($i=2,4,8$) found in Bovet et al. (2022) correspond to distinct topological characteristics of the sequences of partitions, as captured by the block structure in the distance between MCbiF Hilbert functions. The Hilbert distances also capture increased time reversibility in $\theta_{\tau_{4}}$ and $\theta_{\tau_{8}}$ due to the larger stability of social groupings over time (see Fig. 14 in Appendix D.3).
@@ 5.3 Application to Real-World Temporal Data
In our final experiment, we apply MCbiF to temporal sequences of partitions computed from real-world contact data of free-ranging house mice. This data captures the changes in the social network structure of the rodents over time (Bovet et al., 2022). Each partition $\theta_{\tau}(t)$ describes mice social groupings for $N=281$ individual mice at week $t\in[1,\dots,9]$ for the nine weeks in the period 28 February–1 May 2017, so that the sequence $\theta_{\tau}$ captures the fine-graining of social groups from winter to spring. A partition sequence is computed at temporal resolution $\tau>0$, where the parameter $\tau$ modulates how fine the temporal community structure is (Fig. 13). See Bovet et al. (2022) and Appendix D.3 for details. We then compare temporal partition sequences $\theta_{\tau_{i}}$ for different resolutions $\tau_{i}$, $i=1,\dots,9$, as given in Bovet et al. (2022). For each $\theta_{\tau_{i}}$, we obtain $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the nerve-based MCbiF, which induces a 50-fold reduction in computation time due to a much lower number of simplices (260 simplices for the nerve-based MCbiF vs. 116,700 for the original MCbiF).
@@
Bovet et al. (2022) identified that the temporal resolutions $\tau_{2}=1\,\mathrm{s}$, $\tau_{4}=60\,\mathrm{s}$ and $\tau_{8}=24\,\mathrm{h}$ lead to robust sequences of partitions. Using the Hilbert distance (i.e., the $L_{2}$-norm on the 0- and 1-dimensional MCbiF Hilbert functions), we find these temporal resolutions to be representative of three distinct temporal regimes characterised by different degrees of non-hierarchy, as measured by $\bar{c}_{0}$ and $\bar{c}_{1}$ (Fig. 3). In particular, high $\bar{c}_{0}$ indicates that mice tend to split off groups over time, and high $\bar{c}_{1}$ indicates that mice meet in overlapping subgroups but never jointly in one nest box. Note that $\theta_{\tau_{2}}$ has a strong non-hierarchical structure because the large-scale mice social clusters get disrupted in the transition to spring. In contrast, $\theta_{\tau_{8}}$ is more hierarchical as it captures the underlying stable social groups revealed by the higher temporal resolution. Yet, $\theta_{\tau_{4}}$ has the strongest hierarchy as indicated by a lower $\bar{c}_{0}$ and an absence of 1-conflicts ($\bar{c}_{1}=0$) and thus corresponds to a sweet spot in hierarchical organisation between low and high temporal resolutions. These findings also underpin the optimised Sankey diagrams in Figure 3a. While the Sankey diagram for $\theta_{\tau_{4}}$ can be drawn without any crossings, the optimal Sankey diagram for $\theta_{\tau_{8}}$ has a crossing from week 1 to week 2, which resembles the ‘hourglass’ 1-conflict (Fig. 2d). This is predicted by our MCbiF features: $\mathrm{HF}_{1}(1,2)=1$ for $\theta_{\tau_{8}}$ implying one crossing between the first and second layer of the Sankey diagram that cannot be undone, as per Corollary 19.
@@ 6 Conclusion
We have introduced the MCbiF, a novel bifiltration that encodes the cluster intersection patterns in multiscale, non-hierarchical sequences of partitions. Its stable Hilbert functions $\mathrm{HF}_{k}$ quantify the topological autocorrelation of the partition sequence $\theta$ and measure non-hierarchy in two complementary ways: at dimension $k=0$ it captures the absence of a maximum with respect to the refinement order (0-conflicts), whereas at dimension $k=1$ it captures the emergence of higher-order cluster inconsistencies (1-conflicts). This is summarised by the measures of average 0-conflict $\bar{c}_{0}(\theta)$ and average 1-conflict $\bar{c}_{1}(\theta)$, which are global, history-dependent and sensitive to the ordering of the partitions in $\theta$. The MCbiF extends the 1-parameter MCF defined by Schindler & Barahona (2025) to a 2-parameter filtration, leading to richer algebraic invariants that describe the full topological information in $\theta$. It is important to remark that the MCbiF is independent of the chosen clustering algorithm and can be applied to any (non-hierarchical) sequence of partitions $\theta$. We demonstrate with numerical experiments that the MCbiF Hilbert functions provide topological feature maps for downstream machine learning tasks, which are shown to outperform both baseline features and representation learning on the raw data for regression and classification tasks on non-hierarchical sequences of partitions. Moreover, the grounding of MCbiF features in algebraic topology enhances interpretability, a crucial attribute for explainable AI and applications to real-world data.
@@
Our analysis of the MCbiF MPH is restricted to dimensions 0 and 1 for simplicity. However, the analysis of topological autocorrelation for higher dimensions would allow us to capture more complex higher-order cluster inconsistencies and could be the object of future research. Furthermore, we focused here on Hilbert functions as our topological invariants because of their computational efficiency and analytical simplicity, which facilitates our theoretical analysis. In future work, we plan to use richer feature maps by exploiting the block decomposition of the MCbiF persistence module, which leads to barcodes (Bjerkevik, 2021), or by using multiparameter persistence landscapes (Vipond, 2020). Another future direction is to use MCbiF to evaluate the consistency of assignments in consensus clustering (Strehl & Ghosh, 2002; Vega-Pons & Ruiz-Shulcloper, 2011). It can be shown that the values of the Hilbert function $\mathrm{HF}_{k}(s,t)$ further away from the diagonal ($s=t$) are more robust to permuting the order of partitions in $\theta$ (see Proposition 36 in Appendix B.4). In particular, $\mathrm{HF}_{k}(t_{1},t_{M})$ only depends on the set of distinct partitions in $\theta([t_{1},\infty))$ and is independent of any permutation in their order. Hence, in future work, $\mathrm{HF}_{k}(t_{1},t_{M})$ could be used as an overall measure of consistency in $\theta$ in the vein of consensus clustering. Finally, minimal cycle representatives of the MPH (Li et al., 2021) can be used to localise 1-conflicts in $\theta$, which is of interest to compute conflict-resolving partitions in consensus clustering, or to identify inconsistent assignments in temporal clustering (Liechti & Bonhoeffer, 2020).
@@
Detailed proofs of all theoretical results can be found in Appendix B and extensive documentation of our experiments in Appendix D. An implementation of our method and code to reproduce our numerical experiments is available at: https://github.com/barahona-research-group/MCbiF. The dataset studied in Section 5.3 is publicly available at: https://dataverse.harvard.edu/file.xhtml?fileId=5657692.
@@ Acknowledgments
JS acknowledges support from the EPSRC (PhD studentship through the Department of Mathematics at Imperial College London). MB acknowledges support from EPSRC grant EP/N014529/1 supporting the EPSRC Centre for Mathematics of Precision Healthcare. This work benefited from discussions at the London-Oxford TDA seminar in November 2024. We thank Kevin Michalewicz, Asem Alaa, Christian Schindler and Jacopo Graldi for valuable discussions on computational aspects of this project, Alex Bovet for discussions on the temporal dataset studied in this paper, and Arne Wolf for discussions on the stability of multiparameter persistence modules. We also thank Heather Harrington for helpful discussions and for the opportunity to present work in progress at the Max Planck Institute of Molecular Cell Biology and Genetics Dresden (MPI-CBG) in July 2025.
Author Contributions
JS developed the mathematical theory, conducted all numerical experiments, and authored the initial draft of the paper. MB provided supervision, idea generation and guidance throughout the full research process. Both authors wrote and edited the final manuscript.
@@
Appendix A Summary of Key Theoretical Results
Figure 4: Summary of key theoretical results and their relationships. Double-headed arrows represent equivalences (iff), whereas single-headed arrows represent implications (if).
Appendix B Proofs of Theoretical Results and Additional Details
B.1 Proofs and Details for Section 1
@@ B.2 Proofs and Details for Section 3
Remark 24.
By fixing $s:=t_{1}$ (i.e., the top row in the commutative MCbiF diagram), we recover the 1-parameter Multiscale Clustering Filtration (MCF) defined by Schindler & Barahona (2025). MCF was designed to quantify non-hierarchies in coarse-graining sequences of partitions and thus cannot capture fine-graining. For example, a large cluster $C\in\theta(s^{\prime})$ prevents MCF from detecting cluster assignment conflicts between elements $x,y\in C$ for $t\geq s^{\prime}$, see Section 3.1. In particular, the MCF is not a complete invariant of $\theta$. In contrast, MCbiF is a complete invariant of $\theta$ and encodes the full topological autocorrelation contained in $\theta$ by varying both the starting scale $s$ and the lag tst-s.
Next, we provide formal definitions for algebraic properties of persistence modules, see Botnan & Lesnick (2023) for details.
@@
Let $\mathbf{Vect}$ denote the category of $\bm{k}$-vector spaces for a fixed field $\bm{k}$. For a partially ordered set $P$, a $P$-indexed persistence module is a functor $F:P\mapsto\mathbf{Vect}$. We say that:
@@
    $F$ is called pointwise finite-dimensional if $\dim(F_{\bm{a}})<\infty$ for all $\bm{a}\in P$.
@@
    $F$ is called block-decomposable if it decomposes into blocks $F\bigoplus_{B\in\mathcal{B}(F)}\bm{k}_{B}$ where $\mathcal{B}(F)$ is a multiset of blocks (see Definition 25) that depends on $F$.
We can now provide the proof for Proposition 27, which shows that the MCbiF persistence module (see Fig. 6) is pointwise finite-dimensional, finitely presented and block-decomposable. The proof relies on the equivalent nerve-based construction of the MCbiF (see Proposition 34), and the exactness of the persistence module from which block-decomposability follows (Cochoy & Oudot, 2020).
Proposition 27.
@@
To prove block-decomposability, we use the nerve-based MCbiF $(\tilde{K}^{s,t})_{t_{1}\leq s\leq t}$, which leads to the same persistence module, see Proposition 34. As the module is uniquely defined by its values on a finite grid, we can use Theorem 9.6 by Cochoy & Oudot (2020) that implies block-decomposability if the persistence module is exact. Hence, it suffices to show that for all $t_{1}\leq t\leq t^{\prime}\leq t^{\prime\prime}\leq t^{\prime\prime\prime}$ the diagram
@@
By construction of the MCbiF, $\tilde{K}^{t,t^{\prime\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. Furthermore, $\tilde{K}^{t,t^{\prime\prime}}=\tilde{K}^{t,t^{\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime}}$ and $\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}=\tilde{K}^{t^{\prime},t^{\prime\prime}}\cup\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}$. As $\theta$ is a piecewise-constant function with change points $t_{1}<\dots<t_{M}$, we can assume without loss of generality, $_{k}, t=tt^{\prime}=t_{\ell}$, $^{\prime\prime}=t_{m}, t′′′=t$$^{\prime\prime\prime}=t_{n}$ for change points $t_{k}<t_{\ell}<t_{m}<t_{n}$ of $\theta$ such that $A(k,\ell)\cap A(m,n)=\emptyset$. Hence, $\tilde{K}^{t,t^{\prime}}\cap\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}=\emptyset$ and $\tilde{K}^{t^{\prime},t^{\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cap\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. This means that Eq. equation 7 is a Mayer-Vietoris sequence for all $k\geq 0$, implying exactness (Hatcher, 2002, p. 149) and proving the block decomposability (Cochoy & Oudot, 2020, Theorem 9.6). ∎
@@
B.2.1 Proofs and Details for Section 3.1
Remark 28.
It follows directly from the definitions that a hierarchical sequence $\theta$ is always nested. However, nestedness does not necessarily imply hierarchy, as illustrated by the example in Fig. 2b.
@@
(i) Moreover, $\nexists r\in[s,t]$ such that $x\sim_{r}\sim_{r}y$. In particular, $\nexists r\in[s,t]$ such that $\nexists C^{\prime\prime}\in\theta(r)$ with $C\subseteq C^{\prime\prime}$ and $C\subseteq C^{\prime\prime}$. Hence, $\nexists r\in[s,t]$ such that $\theta(r_{1})\leq\theta(r)$ and $\theta(r_{2})\leq\theta(r)$, implying that the subposet $\theta([s,t])$ has no maximum. This shows that every triangle 0-conflict is also a 0-conflict, proving statement (i). Note that the opposite is not true, as illustrated by the example in Fig. 2b. ∎
Remark 29.